OpenStudy (anonymous):

what decimal is represented by this expanded form 2x10,000+7x1,000+4x100+9x1+3x1/10+8x1/1,000

ganeshie8 (ganeshie8):

\[2\times 10,000+7\times 1,000+4\times 100+9\times 1+3\times 1/10+8\times 1/1,000\] \[2\times 10^4 + 7\times 10^3 + 4\times 10^2 + 0\times 10^1 + 9\times 10^0 + 3/10^1 + 0/10^2 + 8/10^3\] This is the decimal expansion of number which can be expressed more compactly as : \[\large (27409.308)_{10}\]

ganeshie8 (ganeshie8):

when it is clear that the base is 10, you may leave that also and simply write it as \[\large 27409.308\]